3. ltistnown that 65% of colege students like rap muselnagroupof45stdents ind the probability that Less than 27 lke rap music d) More than 33 ke rap music 5 points 4 The following chart compares occupations with salaries Salary Range 0.000 30 40,000 or less 42 to 71 to 100,000 More than 100,000 total 110 115 30 10 35 зо 10 so Total 90 130 410 Find the probability that a person selected from this group at random is cl a mure or someone earning between 41.000 and 70.000 a Accountant give that they earn between 41.000 and 0.000 5 points 5
Answer #1

3) Answer: ---- Date: ----25/11/2018

Since n is large .We use normal approximation to binomial, if the necessary conditions are satisfied i.e. np> 5,nq> 5 np = 45*0.65 = 29.25 > 5 nq-45*0.35-15.75 > 5 Let X be the number of college student like rap music Mean μ = np-45 * 0.65 = 29.25 Standard deviation-ơ vnpq V45 * 0.65 * 0.35-3.1996 X N( 29.25,3.1996) a) P(X 31)-P(30.5< X< 31.5)- P 30.5 - 29.25 3.2 31.5 -29.25 3.2 P(0.39 < Z < 0.70) = 0.2580-0.1517 = 0.1063

4) Answer: ----